# Controlling Large Language Model Agents with Entropic Activation Steering

### v1

@@ 6.2 Understanding steering vectors
Effectiveness of steering vectors at different target layers.   EAST requires a choice of the layer in the LLM that will be used during its two phases, with an impact on both the computation of the steering vector, and on the application of the vector during the interactions of the agent with the environment.
We show in Figure 7 that, regardless of the choice for the multiplier $\beta$, the layers at which EAST’s intervention is most effective sit in the middle of the LLM, with a peak at the 16th layer, which we used in the rest of our experiments. This is in line with previous work on interpreting the representations of pretrained LLMs outside of the agentic setting, which found that the representation of abstract concepts such as sycophancy and refusal resides in layers roughly in the middle of the LLM [21].
@@
7 Discussion
7.1 Limitations and Future Work
The current version of EAST is designed for environments with discrete actions, in which it is easy to estimate the entropy of the action distribution. Future work can generalize the method to settings in which actions are open-ended, such as in software engineering [6] or tool-use [22] applications.
We have shown that EAST allows to control the amount of exploration an LLM agent exhibits. This does not automatically prescribe how much an agent should explore in a given task, or how its exploration behavior should evolve over time. However, we believe that EAST can be used as a building block to steer an agent’s exploration in future LLM-based systems. We also believe that EAST’s demonstration that LLMs explicitly represent uncertainty about actions can inform the design of such systems.
7.2 Conclusions
In this paper, we studied how in-context LLM agents behave in sequential decision-making tasks and how they represent uncertainty. After having established that they tend to be overconfident about their decisions, we introduced Entropic Activation Steering (EAST), a method for influencing their behavior and actions. We illustrated how token-level sampling and action generation interact, and demonstrated that EAST can increase the entropy of an LLM agent’s action distribution and alleviate its overconfidence, well-beyond what is achievable by increasing increasing the sampling temperature at the token level. In addition, we have shown that EAST is able to modify the subjective uncertainty of an LLM agent, influencing its thoughts towards more uncertain and explorative attitudes.
As designers of agentic LLM-based systems, it is paramount for us to be able to interpret how they make decisions and to steer them towards more desirable behaviors. EAST advances our understanding of the representation that in-context LLM agents have about their uncertainty over decisions, and our ability to control it. Considering that uncertainty over one’s actions is such a fundamental aspect of successful decision-making, we believe our work to be a promising step in the development of interpretable and steerable in-context LLM agents.
@@ A.1.2 Language Model Assets
We conduct experiments on Mixtral-8x7b model [5], available at this link, and the DBRX model [2] available here. Mixtral is released under the Apache 2.0 license, and DBRX is released under the Databricks Open Model License.
@@ A.1.4 Computation of Relative Frequency
A.2 Additional Experimental Results
DBRX experiments.
While the focus of our study is the Mixtral-8x7b model, we conduct additional experiments on the DBRX open LLM. We repeat experiments detailed in Sections 4.1 and 6.1 using this model. The results pictured in Figure 10 show that this model behaves similarly to Mixral-8x7b, both in its default strategies on the bandit task and its response to the EAST intervention.
A.3 Prompts and Example Interactions

### v2

@@ 6.2 Understanding steering vectors
Effectiveness of steering vectors at different target layers.   EAST requires a choice of the layer in the LLM that will be used during its two phases, with an impact on both the computation of the steering vector, and on the application of the vector during the interactions of the agent with the environment. We show in Figure 7 that, regardless of the choice for the multiplier $\beta$, the layers at which EAST’s intervention is most effective sit in the middle of the LLM, with a peak at the 16th layer, which we used in the rest of our experiments. This is in line with previous work on interpreting the representations of pretrained LLMs outside of the agentic setting, which found that the representation of abstract concepts such as sycophancy and refusal resides in layers roughly in the middle of the LLM (Rimsky et al., 2023).
@@
6.3 Generalizing EAST across tasks and LLMs
Effectiveness of EAST on other LLMs.   As mentioned at the beginning of the section, we employed a Mixtral-8x7b model in most of our experiments, since it provides a good tradeoff between inference speed and performance. To demonstrate the generality of EAST, we conduct additional experiments on the DBRX open LLM (Databricks, 2024). We repeat experiments detailed in Section 4.1 and Section 6.1 using this model. The results pictured in Figure 10 show that this model behaves similarly to Mixral-8x7b, both in its default strategies on the bandit tasks and its response to the EAST intervention. This demonstrates that, despite the specific information encoded in an LLM’s representation depends on its training data and the exact training procedure that was used to train it, EAST can correctly identify the appropriate direction in the representation space and intervene on the LLM agent’s behavior.
7 Conclusions
In this paper, we studied how in-context LLM agents behave in sequential decision-making tasks and how they represent uncertainty over actions. After having established that they tend to be overconfident about their decisions, we introduced Entropic Activation Steering (EAST), a method for influencing their exploration. We illustrated how token-level sampling and action generation interact, and demonstrated that EAST can increase the entropy of an LLM agent’s action distribution and alleviate its overconfidence, well beyond what is achievable by increasing the sampling temperature at the token level. In addition, we have shown that EAST can modify the subjective uncertainty of an LLM agent, influencing its thoughts towards more uncertain and explorative attitudes.
We believe that EAST can be used as a building block to steer an agent’s exploration in future LLM-based systems and that EAST’s demonstration that LLMs explicitly represent uncertainty about actions can inform the design of such systems. As designers of agentic LLM-based systems, it is paramount for us to be able to interpret how they make decisions and to steer them towards more desirable behaviors. EAST advances our understanding of the representation that in-context LLM agents have about their uncertainty over decisions, and our ability to control it. Considering that uncertainty over one’s actions is a fundamental aspect of successful decision-making, we believe our work to be a promising step in the development of interpretable and steerable in-context LLM agents.
@@ A.1.2 Language Model Assets
We conduct experiments on Mixtral-8x7b model Jiang et al. (2024), available at this link, and the DBRX model Databricks (2024) available here. Mixtral is released under the Apache 2.0 license, and DBRX is released under the Databricks Open Model License.
@@ A.1.4 Computation of Relative Frequency
A.2 Prompts and Example Interactions